OpenStudy (anonymous):

The function g(x) = x2 – 10x + 24 is graphed on a coordinate plane. Where will the function cross the x-axis? (–6, 0) and (–4, 0) (–2, 0) and (–12, 0) (4, 0) and (6, 0) (12, 0) and (2, 0)

OpenStudy (anonymous):

@AlexandervonHumboldt2 @triciaal @Michele_Laino @Starbucks

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Function will cross x axis when y = 0. Put y = 0 and find x

OpenStudy (alexandervonhumboldt2):

guess C

OpenStudy (anonymous):

@AlexandervonHumboldt2 Your Unsure

OpenStudy (michele_laino):

at point x, such that: \[x ^{2}-10x+24=0\]

OpenStudy (alexandervonhumboldt2):

yeah i'm bad with functions

OpenStudy (anonymous):

y = x^2 - 10x + 24 sum of factors = -10 product of factors = 24 a+b = -10 a*b = 24 find a and b

OpenStudy (triciaal):

are we using -6 and -4 to get -10 when added and +24 when multiplied?

OpenStudy (michele_laino):

Sorry I think x=6, and x=4

OpenStudy (triciaal):

@Michele_Laino correct

OpenStudy (anonymous):

Answer Choice

OpenStudy (michele_laino):

because I haveto applid this formual: \[x=\frac{ -b \pm \sqrt{b ^{2}-4*a*c} }{ 2a }\]

OpenStudy (michele_laino):

oops because I have applied...

OpenStudy (anonymous):

@Michele_Laino is correct.

OpenStudy (michele_laino):

thanks! @triciaal

OpenStudy (anonymous):

A B C Or D

OpenStudy (triciaal):

i just answered some more questions for this poster. my point is to get the correct factors to use to set the function = 0

OpenStudy (michele_laino):

thanks! @sangya21

OpenStudy (anonymous):

@Michele_Laino A B C Or D

OpenStudy (triciaal):